OpenStudy (anonymous):

i know this ones supposed to be really easy but i want to make sure i did it right.Convert to rectangular form,r=-3sin(theta)

OpenStudy (anonymous):

r^2 = 9sin^2 (theta)

OpenStudy (anonymous):

so x^2 + y^2 = 9y^2/ (x^2 + y^2)

OpenStudy (amistre64):

r=-3sin(theta) r^2 = -3 rsin(t) x^2 + y^2 = -3y

OpenStudy (anonymous):

is the answer x^2+y^2-3y=0?

OpenStudy (amistre64):

x^2 + y^2 +3y = 0 is what I get, if it done right

OpenStudy (anonymous):

ohh +3y..righttt..got it! thanks again!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

x^2 + y^2 + 3y = 0 means r^2 = -3rsin(theta)

OpenStudy (anonymous):

ur question says -3sin(theta) not -3rsin(theta)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

sooo...wat do i do

OpenStudy (anonymous):

look at wht i wrote and simplify it

OpenStudy (anonymous):

no its right

OpenStudy (anonymous):

its fyn leave it

OpenStudy (anonymous):

ohk:)